18 December 2008 times are MST, UT = MST + 7h

morning The new F972N20 filter in the ring does not fit into the hole in the filter wheel 2. For clarification, the wheel closest to the cryostat is wheel 2 and now has 3 filters & 2 empty slots: empty, r-SLOAN, empty, empty, g-SLOAN. Wheel 1 is closest to the primary and has 4 filters & 1 empty slot: empty, U-BESSEL, V-BESSEL, SDT_Uspec, B-BESSEL.

~14:00 Took 100 biases with telescope at horizon. No problem.

14:30-14:50 Scheduled power outages to check switch. Didn't switch to generator automatically.

15:05 Sent StoneC_UBVVRI.ob with 3 times exposure time to activeate trackers and see how camera runs. This is the OB being used when we first had camera communication problems last night. Telescope is locked out at horizon and telescope and guiding boxes are unticked. Sequence starts with lbcb.20081217.220627.fits. Dome temperature is 2.13C - about same as last night (when it was ~1.86 C).

14:30-16:45 Numerous planned power outages.

afternoon Doug & Aaron went up to the weather station to investigate why temperature and humidity readings were not being updated. They found a cable that was disconnected - the connection should be sturdier to withstand the weather.

17:00 ... UT date now 20081218 Fernando, Andrea, Vincenzo loading new firmware for Red controller. Also testing a way to reset the camera without powering off/on the PCs on red; will be implemented for both Red and Blue.

18:10 Turning on LBC-Blue. Plan to run through a set of 'fake' observations to reproduce the errors of last night.

20:15 Telescope is up and ready. We are running a series of test images at a range of elevations to see if and when the communication problem, if elevation dependent - such as a stressed fiber, will arise. Using fastOB to move the telescope to various elevations. Note that the AZ-EL system can be enabled and it looks like you can specify coordinates AZ-EL, but these are interpreted as RA-DEC by the User Interface.

20:30 Taking data at elevations: 5, 27, 57, 66, 77 degrees.

21:16 Back at 27 deg elevation - so far no communication errors.

22:00 Loaded a fastOB to reproduce the position when it failed last night on the StoneC OB around 00:30am: HA=2h, elev=48deg. Played this OB, added to this OB 5 dithers and 3 filters. So far, no communication failures. This OB was to obtain 3 U-BESSEL exposures, 1 B-BESSEL one and 1 V-BESSEL one at each of 5 dither positions.

22:41 SPC-A200 status register read failed. Elev=50 deg, HA=+02:16. The error occured during the OB, after the 3rd U- BESSEL exposure was complete. It failed before obtaining a B-BESSEL image. Vincenzo powering off and on the camera. Did not come up - camera & shutters failed to load LCA on SPC-A200 error.

2008/12/18 05:53:47.443000 E B CAMERA CCDCTRL load LCA on SPC-A200 error (SPC-A200 status register read fail) [ccdctrl.c:824]

2008/12/18 05:53:47.437147 E B CAMERA CCD Controller error, errcode:8 [src/camera/camera.c:552]

Second attempt at cycling power worked.

23:05 Send another OB, again at HA=2, elev=50. Failed with SPC-A200 register read error on 3rd U-BESSEL exposure. Vincenzo sent another OB to move it to elev=80, but it failed immediately after starting this OB, even before it reached the target.

23:11 Slew to elev=72 degrees and turn on camera. It failed. At 06:15:46 UT in lbc.log there is the message "SPC-A200 status register read fail" (code:38) controller error [ccdctrl.c:2207].

00:03 Trying an OB which does not dither but takes 3 images each at B, V and U. Failed at 3rd exposure at B. Took 2 (070541 and 070720). Turned off/on camera but failed on startup. Elevation = 82 degrees. Trying again. Success.

00:23 Playing an OB with 5 dithers but only 1 filter, B. Elevation = 85 degrees. Completed with no problem but note that the dither offset was 0".

00:33 Playing an OB with 5 dithers (again 0" dither), 1 filter (B, again) but, this time, 2 images per dither position. Same elevation (85 deg). Starts with 073443. Worked.

00:52 Playing a similarly structured OB as the previous one, but with the V instead of B filter.

01:12 Playing an OB which does 5 dithers (0") and at each position, an image at B and V. Starts with 081204. Elevation = 77 degrees. Completed without error.

01:30 Try the same OB as before but with U and B in place of B and V. Starts with 082917. Realized dither offset was 0", so stopped OB and set dither offset to 40". Also reset filters to B and V to try 01:12 experiment again. This OB started with 083454, but it failed right after and image 083454 does not appear in the Repository or in the lbcarchive. Power cycled - on first try camera did not come up. Second try OK.

01:37 Going back to repeat the OB with 5 dithers and 1 filter since the previous times we had 0" dither offset. Running an OB with 5 dithers, B filter, 2 exposures per dither. Elevation = 70 degrees. Completed successfully.

02:00 Trying same with V filter: 5 dithers, 2 exposures per dither. Elevation = 67 degrees. Starts with 090411. After 090550 completed and the timer was reading 60sec/60sec, the User Interface still reported Acquisition Status as "exposing". Finally it ended with an RPC Client/Host communication error at 09:09:27 UT (in lbc.log). The User Interface never reported that this image, 090550, was readout and it never appeared in the lbcarchive or Repository.

Power cycled LBC, but camera & shutter did not come up on the first try (error while turning off) or on second (while turning on we had a "load LCA on SPC-A200 error" ... 09:17:07 in lbc.log). Came up on 3rd attempt.

02:25 Try an OB with 5 dithers and 4 filters per position: U-BESSEL (20s), g-SLOAN (20s), r-SLOAN (20s) and B-BESSEL (20s). This is to avoid V-BESSEL which we are suspicious causes problems. Started with 092604 and went through all filters at the first dither position and all filters at the second but it did not save the B-BESSEL image. It crashed on image 093406 which was at U-BESSEL (the first filter of 3rd dither position). User Interface shows Acquisition Status is "exposure" and yet Exposure time is 20/20. User Interface hung here - Vincenzo gave an lbckill/lbcstart, followed by turning on LBC - RPC Client/Host error on turning on. Turn off/on again. Again an error (load LCA on SPC-A200 error at 09:54:19 in lbc.log). Bringing telescope to zenith. Turn on now.

03:00 Now a 5-filter (g,r,B,U,V) OB with 2 dithers but 0" dither offset. Started with 100537 and crashed on 2nd dither with SPC-A200 register read fail error. The last image saved was the r-SLOAN image, 2nd one in the filter sequence, 101248.

03:25 Going to shut down. The communication errors seem to arise when we run an OB which combines dithers and filter changes. These seem to happen on the 2nd or 3rd dither position, or at least after the 1st position. And tonight we had both the SPC-A200 and RPC errors. SPC-A200 errors started at 05:41:38 UT (22:41 local time) and RPC errors only started after 09:07:27 UT (02:00 local time).

Telescope parked at horizon. Mark and Doug plan to look at the fit of the new filter again.
Topic revision: r6 - 18 Dec 2008, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback